OpenStudy (anonymous):

How do I solve the following: 2 1/2 + 1 1/4 - 2/3 x 6

OpenStudy (january123):

Well, i googled it and got -0.25, but idk if it's right tho...Tell me what u think

OpenStudy (january123):

@sheelo_mughal

OpenStudy (anonymous):

\[2\frac{ 1 }{ 2 }+1\frac{ 1 }{ 4 }-\frac{ 2 }{ 3 }\times6\]

OpenStudy (anonymous):

is this correct question

OpenStudy (january123):

i believe so, don't you change 2 1\2

OpenStudy (anonymous):

\[\frac{ 5 }{ 2 }+\frac{ 5 }{ 4 }-\frac{ 2 }{ 3 }\times6\]

OpenStudy (january123):

okay, @calliope11p why don't you do the problem that she has broken down for you and tell us what you get

OpenStudy (january123):

what is 5+5?

OpenStudy (anonymous):

y r u asking me

OpenStudy (anonymous):

u didnt posted the question

OpenStudy (anonymous):

if @calliope11p wants to ask then no prob

OpenStudy (anonymous):

I came up with the answer of -1/4

OpenStudy (anonymous):

But I am unsure if that is correct... I worked it out, following BODMAS as follows:

OpenStudy (january123):

not you @sheelo_mughal

OpenStudy (january123):

you mean PEMDAS?

OpenStudy (january123):